OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y= -12 A. y =-(11/9) x- 4/3; slope: 11/9; y intercept: 4/3 B. y =-(11/9) x+ 4/3; slope: 11/9; y intercept: 4/3 C. y =11/9 x- 4/3; slope: 11/9; y intercept:-(4/3) D. y =11/9 x+ 4/3; slope: 4/3; y intercept: 11/9

hartnn (hartnn):

to isolate y, add 11x on both sides, what u get?

hartnn (hartnn):

u get 9y=11x-12 then divide by 9,to get y=11/9x-12/9 and -12/9=-4/3 so whats the answer?

OpenStudy (anonymous):

C? I don't really understand this i'm sorry.

hartnn (hartnn):

yup C